OpenStudy (anonymous):

c + (4 – 3c) – 2 = 0?

OpenStudy (austinl):

Solve for c?

OpenStudy (anonymous):

yes

OpenStudy (austinl):

Ok, so you have to distribute the positive to get \[c+(4-3x)-2=0\] \[c+4-3x-2=0\] Then you combine like terms.

OpenStudy (anonymous):

I'm sorry, i just don't understand, but thanks for trying!

OpenStudy (austinl):

Ok, combine terms with c. And combine numerical terms. Can you do that?

OpenStudy (anonymous):

yea i did, it didn't come out right lol. i know the steps but i want to know how you can go from c + (4 – 3c) – 2 = 0 to -2c + 2=0 . that is what I'm confused about.

OpenStudy (marissalovescats):

Then you just solve for c! :)

OpenStudy (anonymous):

so when you distributec all your doing is adding c to everything, and taking out the parenthesis

OpenStudy (anonymous):

for example if i give you 2+(4-3) your going to add 2+4 and 2 =3 and you will get 6+5, so you do the same with your equations all you have to do it take out your parenthesis and solve it so austin gave you that now you combine like terms , you subtract 4 from 2